OpenStudy (anonymous):

Solve The System Please Help!! x-y+z=9 x+y+z=5 x+y-z=-3 A) nothing B) (3, -2, 4) C) (3, 4, -2) D) (4, 3, -2)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

x-y+z=9------1 x+y+z=5------2 x+y-z=-3------3 Adding 1 &2 we find 2x+2z=14 similarly subtracting 3 from 2 we find 2z=8 i.e z=4 so 2x+2z=14 i.e. 2x+8=14 i.e. 2x=6 i.e. x= 3 From --2 we have x+y+z=5------2 i.e. 3+y+4=5 i.e. y+7= 5 i.e. y=-2 thus option B) (3, -2, 4) is correct ans